### 26 Mar 2026 v1

# Circular Future of Yukawa Unification in Supersymmetric Pati-Salam Model: Stop on Neutralino

Circular Future of Yukawa Unification in
Supersymmetric Pati-Salam Model:
Stop on Neutralino
Ali Çiçi [^1] E-mail: ali.cici@cern.ch,a,b, Büşra Niş [^2] E-mail:501507008@ogr.uludag.edu.tr,c and Cem Salih $\ddot{\rm U}$n [^3] E-mail: cemsalihun@uludag.edu.tr,c
[…]
bDepartment of Physics, Faculty of Engineering and Natural Sciences,
Bursa Technical University, TR16310, Bursa, Türkiye
[…]
We explore the low scale implications of SUSY models based on the gauge symmetry $SU(4)_{C}\times SU(2)_{L}\times SU(2)_{R}$. We include the non-holomorphic terms arises from the perturbations on $D$-branes, which also break SUSY. These terms significantly change the implications of Yukawa unification, since they are directly included in the threshold contributions to Yukawa couplings. With these contributions, YU can be compatible with low fine-tuning and yet heavy Higgsino-like LSP, but these solutions receive strong negative impact from the dark matter observations. Besides, in contrast to the earlier results, the NH contributions accommodate heavy gluino masses from about 2.2 to 10 TeV compatible with Yukawa unification. These gluinos can be probed up to about 2.5 in the collider experiments of high luminosity, and to about 6 TeV in future experiments of 100 TeV center of mass energy. In contrast to the gluino mass scales, the NH contributions can drive the squarks of the third family to the light masses. We observe that the Yukawa unification solutions can be compatible with the sbottom mass of about 1.5 TeV, and they are more likely to be tested soon. The stop can also be as light as about 1.5 TeV, and it can be nearly degenerate with the LSP neutralino. Even though these solutions are beyond the sensitivity of the current collider analyses, they can be subjected to be tested in future experiments such as those of Future Circular Collider proposal. We project the current analyses for stop-neutralino degenerate solutions to the Future Circular Collider experiments, and realize that the signal processes with semi-leptonic final states can yield about $5\sigma$ significance with appropriate cuts on the kinematic variables. Such a large significance can be realized when $\mathcal{L}\simeq 97$ fb-1. We also present a slight improvement by performing BDT analyses, which can yield a similar significance with $\mathcal{L}\simeq 54$ fb-1.
1 Introduction
Despite excessive studies on supersymmetry (SUSY) and the lack of direct signals from the ongoing experiments, it is still one of the appealing candidates for new physics models beyond the Standard Model (BSM). SUSY models provide dynamically interfering new particles in experimentally testable processes such as pair productions in collider experiments ATLAS:2023afl ; CMS:2021eha , rare decays of $B$-meson Boubaa:2022xsk , muon anomalous magnetic moment (muon g2g-2) Moroi:1995yh ; Martin:2001st ; Giudice:2012pf etc. Indeed, the abundance of testable contributions might be the reason that the null experimental results strike SUSY more than any other BSM models; however, its interference still makes the phenomenological analyses insightful for direct and indirect tests of SUSY (for a recent review, see Constantin:2025mex ). One of the main motivations behind the SUSY models is the solution to the hierarchy problem, and if one assumes soft SUSY breaking (SSB), the supersymmetric models can still provide a cancellation in quadratic divergent contributions to the Higgs boson mass Gildener:1976ai ; Gildener:1979dd ; Weinberg:1978ym ; Susskind:1978ms ; Veltman:1980mj ; Susskind:1982mw ; Dvali:1995qt . Besides, the 125 GeV Higgs boson mass destabilizes the vacuum at about $10^{10}$ GeV Elias-Miro:2011sqh , while the stability of the Higgs potential can be ensured in SUSY models Degrassi:2012ry ; Bezrukov:2012sa ; Buttazzo:2013uya ; Kitahara:2013lfa ; Carena:2012mw . In addition to such strong motivations, the gauge couplings of the Standard Model (SM) can unify at a grand unification scale ($M_{{\rm GUT}}$). In this context, one can construct supersymmetric grand unified theories (SUSY GUTs) and explore possible high scale origins of the particle phenomena observed or expected at the low scale experiments.
Among SUSY GUTs, especially those based on $SO(10)$ provide a special scheme of unification. In addition to the gauge couplings, one can also unify all the matter fields of a family in its 16-Dimensional spinor representation. Unifying all the fields in a single representation also allows to unify all the Yukawa interactions in its minimal construction (hereafter refers to Yukawa Unification - YU) Ananthanarayan:1991xp ; Ananthanarayan:1992cd . The first thing to note is that YU fails to be consistent with the observed masses of the fermions from the first two families Langacker:1980js ; Babu:1992ia . This conflict can be recovered by extending the Higgs sector at $M_{{\rm GUT}}$ by including fields from different $SO(10)$ representations (for a recent reviews see, for instance, Antusch:2019avd ; Saad:2025cfb ). In such cases, the MSSM Higgs fields are, in general, happen to be a superposition of these fields and hence YU is broken. However, YU can be maintained for the third family if one assumes that if the MSSM Higgs doublets are solely resided together in 10-Dimensional ($10_{H}$) representation of $SO(10)$, and the third family fermions acquire their masses only from the vacuum expectation values (VEVs) in $10_{H}$ AdeelAjaib:2013dnf ; Joshipura:2012sr . Once YU can be maintained for the third family, it also yields a significant impact on the low scale implications. This impact arises from the sensitivity to the threshold corrections to the yukawa couplings at the scale where SUSY particles decouple from the spectrum ($M_{{\rm SUSY}}\equiv\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}$). YU requires large negative threshold corrections to $y_{b}$ Gogoladze:2010fu , which can directly shape the masses of stops, sbottom, gluinos and Higgsinos as well as $\tan\beta$ and SSB trilinear scalar couplings Pierce:1996zz . Even though its impacts on the low scale implications have been extensively explored over years, one can still find recent studies Antusch:2025rbp ; Gao:2025gyo ; Hussain:2025bxp ; Li:2025ypb ; Shafi:2023sqa ; Ahmed:2022ibc ; Aboubrahim:2021phn .
Within these studies, the supersymmetric high scale models based on $SU(4)_{C}\times SU(2)_{L}\times SU(2)_{R}$ gauge symmetry (4-2-2 for short) Pati:1974yy ; Lazarides:1980tg ; Kibble:1982ae form an interesting class. It is the maximal subgroup of $SO(10)$ and it emerges if $SO(10)$ breaks through the VEVs in $54_{H}$ and/or $210_{H}$ representations Kibble:1982dd ; Lazarides:1985my ; Babu:2016bmy . The models in this class can also be related to the string theory through the intersecting $D$-branes Mansha:2025gvr ; Leontaris:2025yuh ; Li:2022cqk ; Florakis:2021bws ; Antoniadis:1990hb . The 4-2-2 model preserves most of the salient features of $SO(10)$ such as the conservation of BLB-L symmetry, a kind of matter unification (the matter fields are divided into two representations regarding to their handedness) and YU in its minimal construction. Even though it does not have to be classified in GUTs, if it also breaks into the SM gauge group near $M_{{\rm GUT}}$, the gauge coupling unification can be maintained approximately. If breaking 4-2-2 happens through VEVs in $126$ and $\overline{126}$, then a $Z_{2}$ symmetry is preserved, which acts as the matter parity such that the lightest supersymmetric particle (LSP) happens to be stable Kibble:1982dd . Despite yielding rich phenomena Gogoladze:2009bn ; Shafi:2023ksr ; Gomez:2020gav ; Shafi:2015lfa ; Karagiannakis:2012sv , the YU condition sweeps away most of the implications since it leads to rather heavy mass spectra for the supersymmetric scalars Gogoladze:2011db ; Blazek:2003wz ; Raza:2014upa . Such a spectrum obviously necessitates much stronger colliders with quite large center of mass energy such as those proposed in Future Circular Collider (FCC) FCC:2025lpp ; FCC:2025uan ; FCC:2025jtd . One of our previous analyses has shown that FCC can probe the stops through proton collisions up to about 2.5 TeV at $95\%$ Confidence Level (CL), and about 3 TeV at $68\%$ CL Altin:2020qmu . In contrast to the heavy scalar spectrum, YU in the minimal setup for these models yield relatively light gluinos ($m_{\tilde{g}}\lesssim 1.5$ TeV), and most of these solutions are excluded by the gluino analyses ($m_{\tilde{g}}\gtrsim 1.4$ TeV when it is next to LSP (NLSP) ATLAS:2021yij ; ATLAS:2021twp ; ATLAS:2020syg ).
In our work we revisit this class of Yukawa unified models in the presence of non-holomorphic (NH) SSB terms. These terms are strictly forbidden by the holomorphy condition of SUSY, but they can emerge after the SUSY breaking Inoue:1982pi ; Hall:1990ac ; Bagger:1993ji ; Bagger:1995ay ; Martin:1999hc ; Jack:1999ud ; Jack:1999fa ; Haber:2007dj . The presence of these terms can be a sign for the “hard” SUSY breaking which may bring the hierarchy problem back. However, if there is no singlet field under the SM gauge group, such NH terms can be classified in the SSB terms as well Inoue:1982pi ; Martin:1999hc ; Jack:1999ud ; Jack:1999fa . Despite the abundance of earlier studies on YU in this class of models, almost none of them does consider the contributions from the NH terms. The possible reason behind neglecting them might be that the NH terms are induced at higher loops; hence, they are suppressed by $1/M$ compared to the holomorphic SSB terms Bagger:1995ay ; Martin:1999hc ; Haber:2007dj . Therefore, such terms are negligibly small in gravity mediated SUSY breaking models ($M\simeq M_{{\rm GUT}}$). On the other hand, the NH terms can be induced independently on the SUSY breaking scale within the string theory through the perturbations on $D$-branes deWit:1996kc ; Bellisai:1997ck ; Dine:1997nq ; Gonzalez-Rey:1998xrq ; Buchbinder:1998qd ; McGuirk:2012sb . The NH terms and their impact on the low scale implications have recently received attention, and the earlier studies Chattopadhyay:2017qvh ; Ali:2021kxa ; Chakraborty:2019wav ; Un:2014afa ; Rehman:2022ydc ; Israr:2025cfd ; Israr:2024ubp ; Rehman:2025djc ; Rehman:2024tdr ; Un:2023wws ; Hicyilmaz:2021onw ; Nis:2025fxc have shown that they can drastically change the low scale implications of SUSY models.
The recent analyses on the NH contributions are also promising to explore their impact on YU and its low scale implications. The renormalization group equations (RGEs) for the Yukawa couplings may not be affected from the NH terms directly, but YU can still be sensitive to these terms through the threshold corrections. In this context, one can realize totally different regions for YU in the fundamental parameter space, which also leads to drastically different implications for the low scale observations. Their effects on the SUSY mass spectra can be of a special interest for the direct detection/probe the SUSY particles. YU, in general, favors the heavy mass spectra for strongly interacting SUSY scalars, which can even be beyond the reach of the future colliders. In this case, if the NH terms can reduce their masses considerably, they can be subjected into interesting probes in the collider experiments, as well. In our work, we also discuss possible collider probes after discussing the phenomenological impacts from the NH terms. The rest of the paper is organized as follows: We first discuss the NH threshold corrections to Yukawa couplings which are crucial for YU in Section 2. We also briefly discuss if naturalness of SUSY models can be satisfied. Our numerical and collider analyses together with the experimental constraints are summarized in Section 3. We start the discussion of the NH contributions by comparing their results with the holomorphic cases, and possible probes in Section 4. Section 5 spare discussions on the mass spectra for the SUSY particles with a special emphasize on the stop and LSP neutralino mass scales. After we go into detailed detector analyses to probe stops, we summarize and conclude our findings in Section 6.
2 NH Contributions to the Yukawa Unification
One of the first observations about the impact of NH terms is that excluded solutions in the parameter space of models can become consistent again with the current experimental results after the NH contributions are taken into account Un:2014afa ; Nis:2025fxc . Even though they can widely open the parameter space to be tested experimentally, we consider rather direct NH contributions to the Yukawa couplings that accommodate YU solutions within the models in 4-2-2 class. As mentioned before, YU requires $y_{b}$ to receive large negative threshold contributions, which are diagrammatically represented in Figure 1.
Figure 1: The threshold corrections to the b-quark Yukawa coupling.
$M_{3}$ in the left diagram represents the SSB mass term for the gluino as $m_{\tilde{g}}=|M_{3}|$. The $\mu$-term appears in the diagrams from the mixing of the SUSY scalars in the left diagram, and from the Higgsino mass in the right diagram. Note that the diagrams display only the holomorphic contributions, which can be calculated as Hall:1993gn :
$\delta_{y_{b}}^{H}\approx\dfrac{g_{3}^{2}}{12\pi^{2}}\dfrac{\mu M_{3}\tan\beta}{m_{\tilde{b}}^{2}}+\dfrac{y_{t}^{2}}{32\pi^{2}}\dfrac{\mu A_{t}\tan\beta}{m_{\tilde{t}}^{2}}\penalty 10000\ \penalty 10000\ ,$ (1)
where $m_{\tilde{b}}^{2}=m_{\tilde{b}_{L}}m_{\tilde{b}_{R}}$ and $m_{\tilde{t}}^{2}=m_{\tilde{t}_{L}}m_{\tilde{t}_{R}}$. The first term provides rather positive contributions when $\mu$ and $M_{3}$ are set to be positive in numerical calculations. Thus, it should be suppressed to realize the YU solutions, and such a suppression can be realized with light gluinos ($M_{3}\lesssim 1.2$ TeV) and heavy sbottoms ($m_{\tilde{b}}\gtrsim 5$ TeV) Gogoladze:2009ug . One can also consider lower values for $\mu$ to minimize the positive contributions, but it is included also in the second term which can be accounted for the desired negative contributions. Since the second term is the only source for the negative contributions, YU restricts the trilinear SSB scalar interaction term ($A_{t}$) to be negative. However, if one sets $\mu$ to low values, then the required negative contributions necessitate so large $A_{t}$ that leads to color/charge breaking (CCB) vacua Ellwanger:1999bv ; Camargo-Molina:2013sta . Indeed, previous studies have shown that YU favors large $\mu$ values along with negative $A_{t}$, and the positive contributions from the first term are suppressed rather by the light gluino and heavy sbottom.
The inclusion of the NH terms into the diagrams given in Figure 1 can be seen from the following NH SSB Lagrangian Jack:1999ud ; Jack:1999fa :
$\mathcal{L}^{{\rm NH}}_{{\rm soft}}=-\mu^{\prime}{\tilde{H}_{u}}\cdot{\tilde{H}_{d}}-A_{u}^{\prime}\tilde{Q}\penalty 10000\ {H}_{d}^{\dagger}\tilde{U}-A^{\prime}_{d}\tilde{Q}\penalty 10000\ {H}_{u}^{\dagger}\tilde{D}-A^{\prime}_{e}\tilde{L}\penalty 10000\ {H}_{u}^{\dagger}\tilde{E}-\mbox{h.c.}\penalty 10000\ ,$ (2)
where $\mu^{\prime}$ is the NH term contributing to the Higgsino mass, while $A_{i}^{\prime}$ ($i=u,d,l$) are similar to the holomorphic trilinear scalar coupling. These terms are classified as NH terms such that the SUSY scalars couple to the “wrong” Higgs doublet with $A_{i}^{\prime}$. Note that the family indices are suppressed in Eq.(2), but the third family refers to $i=t,b,\tau$.
The third term in Eq.(2) with $A_{b}^{\prime}$ couples the sbottom to $H_{u}$, and it contributes through the first diagram by $A_{b}^{\prime}y_{b}$. A similar NH contribution can be realized in the second diagram through stop and $H_{u}$ vertex, but such contributions are suppressed by $\tan\beta$, and thus they are negligible in comparison with the contributions from $A_{t}$. On the other hand, the main NH contribution in this diagram appears in the Higgsino mass, which turns to be $m_{\tilde{H}}=\mu+\mu^{\prime}$. Thus, one can summarize the leading NH contributions to $y_{b}$ as follows:
$\delta_{y_{b}}^{{\rm NH}}\approx\dfrac{g_{3}^{2}}{12\pi^{2}}\dfrac{A_{b}^{\prime}M_{3}\tan\beta}{m_{\tilde{b}}^{2}}+\dfrac{y_{t}^{2}}{32\pi^{2}}\dfrac{\mu^{\prime}A_{t}\tan\beta}{m_{\tilde{t}}^{2}}\penalty 10000\ \penalty 10000\ ,$ (3)
and the total threshold contribution becomes $\delta y_{b}=\delta_{y_{b}}^{H}+\delta_{y_{b}}^{NH}$. With the NH terms, the total contribution from the first diagram can turn to be negative even in the cases with $\mu,M_{3}>0$, when $A_{b}^{\prime}$ is negative. Thus one can accommodate YU with heavy gluinos and relatively light sbottoms.
Similarly, negative and large $\mu^{\prime}$ can also lead to desired negative contributions to $\delta_{y_{b}}$ through the second diagram; thus YU can be realized in the regions with low $\mu$ ($\sim\mathcal{O}(100)$ GeV). Hence, the models with YU can be more interesting in naturalness point of view (see, for instance, Baer:2025uzx ). In fine-tuning ($\Delta_{{\rm EW}}$) discussions, we will employ the following definitions Baer:2012mv :
$\Delta_{{\rm EW}}\equiv\dfrac{{\rm Max}(C_{i})}{M_{Z}^{2}/2}$ (4)
with
$C_{i}=\left\{\begin{array}[]{l}C_{H_{d}}=\mid m^{2}_{H_{d}}/(\tan^{2}\beta-1)\mid\penalty 10000\ ,\\ C_{H_{u}}=\mid m^{2}_{H_{u}}\tan^{2}\beta/(\tan^{2}\beta-1)\mid\penalty 10000\ ,\\ C_{\mu}=\mid-\mu^{2}\mid\penalty 10000\ .\end{array}\right$. (5)
Among $C_{i}$ given above, $C_{H_{d}}$ is suppressed by $\tan\beta$; thus, $\Delta_{{\rm EW}}$ is determined mainly by $C_{H_{u}}$ and $C_{\mu}$; and the electroweak symmetry breaking (EWSB) requires $\mid-\mu^{2}\mid\simeq\mid m_{H_{u}}^{2}\mid$. In the absence of NH terms, low $\Delta_{{\rm EW}}$ leads to light masses for Higgsinos, which receive negative and strong impacts from the Planck measurements Planck:2018nkj on relic abundance of dark matter (DM) and direct detection of DM experiments Baer:2024hgq . One can ameliorate this problem with the NH terms that the Higgsinos can be heavier with the contributions from $\mu^{\prime}$. Since $\mu^{\prime}$ does not take part in EWSB, it does not alter the fine-tuning measurement; and hence, one can also fit the heavy Higgsino masses with low $\Delta_{{\rm EW}}$ Un:2023wws .
3 Scanning Procedure, Experimental Constraints and Detector Analyses
In this section, we briefly described the fundamental parameter space of the 4-2-2 class supplemented with the NH terms. The fundamental parameters and their ranges can be summarized as follows:
$\begin{array}[]{lcl}0\leq&m_{0}&\leq 20\penalty 10000\ {\rm TeV},\\ 0\leq&M_{2},M_{3}&\leq 5\penalty 10000\ {\rm TeV},\\ -3\leq&A_{0}/m_{0}&\leq 3,\\ 35\leq&\tan\beta&\leq 60,\\ 0\leq&m_{H_{d}},m_{H_{u}}&\leq 20\penalty 10000\ {\rm TeV},\\ &{\rm sgn}(A^{\prime}_{0}),{\rm sgn}(\mu^{\prime})&=-1,+1\penalty 10000\ ,\end{array}$ (6)
where $m_{0}$ is the universal SSB mass terms for the scalar supersymmetric matter fields, while $M_{2}$ and $M_{3}$ represent the SSB mass terms for $SU(2)_{L}$ and $SU(3)_{c}$ gauginos. The SSB mass for $U(1)_{Y}$ gaugino can be determined by the Pati-Salam relation as
$M_{1}=\dfrac{3}{5}M_{2}+\dfrac{2}{5}M_{3}.$ (7)
$A_{0}$ stands for the universal SSB term for the trilinear scalar interactions, and it is varied with its ratio to $m_{0}$. Its range is restricted as given in Eq.(6) to avoid CCB vacua Ellwanger:1999bv ; Camargo-Molina:2013sta . $\tan\beta$ is the ratio of the VEVs of $H_{u}$ and $H_{d}$ as $v_{u}/v_{d}$, and $m_{H_{u,d}}$ stand for the SSB mass terms for the Higgs doublets.
In our setup the NH terms are assumed to be generated in a scale-independent way through the SUSY breaking perturbations on $D$-branes. Such perturbations induce the NH terms such that $|\mu^{\prime}|=|M_{i}|$ and $|A_{0}^{\prime}|=|A_{0}|$, where $M_{i}$ represent the SSB gaugino terms and we set $|\mu^{\prime}|={\rm max}(M_{2},M_{3})$. Thus, the NH terms do not extend the fundamental parameters of the models. However, we also allow a phase difference between them and their holomorphic partners, and randomly assign them positive and negative signs in our scans.
We perform random scans spanned by the fundamental parameters listed in Eq.(6) by employing the Metropolis-Hastings algorithm Baer:2008jn ; Belanger:2009ti . The random input values are given to SPheno-4.0.4 Porod:2003um ; Goodsell:2014bna numerical calculation package, which was obtained with SARAH Staub:2008uz ; Staub:2015iza . SPheno first runs the 1-loop RGEs for the Yukawa and gauge couplings from $M_{Z}$ to $M_{{\rm GUT}}$, which is determined by the unification condition as $g_{1}=g_{2}\simeq g_{3}$, where $g_{1}$, $g_{2}$ and $g_{3}$ are the gauge coupling for $U(1)_{Y}$, $SU(2)_{L}$ and $SU(3)_{C}$, respectively. Once $M_{{\rm GUT}}$ is determined all the SSB parameters are imposed and SPheno evolves masses and couplings back to $M_{{\rm SUSY}}$ through the 2-loop RGEs, where $M_{{\rm SUSY}}\equiv\sqrt{m_{\tilde{t}_{L}}m_{\tilde{t}_{R}}}$ is the scale at which the supersymmetric particles decouple from the spectrum. In generating data we accept only the solutions which do not yield negative mass-square for the scalar particles and consistent with the radiative EWSB (REWSB), which fixes the bilinear Higgs mixing term $\mu$ with $m_{H_{u}}$, $m_{H_{d}}$ and $\tan\beta$ to be consistent with EWSB. In addition, we require all the solutions to yield one of the MSSM neutralinos to be LSP to involve suitable DM candidate in spectrum. In this context, one can consider LSP neutralino accounted for the DM implications, and so we transfer the SPheno solutions to micrOMEGAs Belanger:2018ccd to add the DM observables in our analyses.
After generating the data, the solutions are subjected to the mass bounds on the supersymmetric particles ParticleDataGroup:2014cgo ; ATLAS:2021twp ; ATLAS:2020syg ; ATLAS:2022rcw and Higgs boson ATLAS:2012yve ; CMS:2012qbp ; CMS:2013btf , constraints from combined results for rare $B$-meson decays CMS:2020rox ; Belle-II:2022hys ; HFLAV:2022esi , and the latest Planck Satellite measurements Planck:2018nkj on the DM relic abundance successively to constrain the LSP neutralino. The list given below summarizes the experimental constraints employed in our analyses:
$\begin{array}[]{l}m_{h}=123-127\penalty 10000\ {\rm GeV}\\ m_{\tilde{g}}\geq 2.1\penalty 10000\ {\rm TeV}\penalty 10000\ (1400\penalty 10000\ {\rm GeV}\penalty 10000\ {\rm if\penalty 10000\ it\penalty 10000\ is\penalty 10000\ NLSP})\\ 1.95\times 10^{-9}\leq{\rm BR}(B_{s}\rightarrow\mu^{+}\mu^{-})\leq 3.43\times 10^{-9}\;(2\sigma)\\ 2.99\times 10^{-4}\leq{\rm BR}(B\rightarrow X_{s}\gamma)\leq 3.87\times 10^{-4}\;(2\sigma)\\ 0.114\leq\Omega_{{\rm CDM}}h^{2}\leq 0.126\;(5\sigma)\penalty 10000\ .\end{array}$ (8)
In addition to these constraints we define the following parameter to measure YU as
$R_{tb\tau}=\dfrac{{\rm max}(y_{t},y_{b},y_{\tau})}{{\rm min}(y_{t},y_{b},y_{\tau})}\penalty 10000\ .$ (9)
The solutions with perfect YU ($y_{t}=y_{b}=y_{\tau}$) corresponds to $R_{tb\tau}=1$. However, we allow $10\%$ deviation to take into account the unknown threshold corrections to the Yukawa couplings at $M_{{\rm GUT}}$ due to the symmetry breaking Hisano:1992jj ; Chkareuli:1998wi . Thus we identify the YU solutions with the condition $R_{tb\tau}\leq 1.1$.
Figure 2: The topology for the signal processes considered in our analyses.
At the end of our analyses we also discuss several benchmark scenarios to consider possible signals which can be probed in the collider experiments. We select some solutions which are consistent with the constraints listed in Eq.(8). In such analyses we consider the stop pair production in which each stop subsequently decays into the LSP neutralino associated with a top quark. The process is diagrammatically shown in Figure 2. Note that we consider the semi-leptonic final states in our analyses due to the detector efficiency in $b$-tagging, which cannot exceed $70\%$ ATLAS:2022miz . While we can keep $b$-tagging efficiency at about $70\%$ in semi-leptonic decays, it would be reduced to about $50\%$ in fully hadronic final states.
In the analyses for the collider implications, we employ MadGraph5_aMC@NLO (v3.6.3) Alwall:2014hca to generate the signal and relevant background events at the parton level. Following the analyses in Ref.CMS:2025ttk , we set the five-flavor-scheme (5FS) and use NNPDF23LO1 NNPDF:2017mvq at this step. The generated events are transferred to Pythia 8.3 Sjostrand:2014zea to complete the final states by parton showering, hadronization and decaying unstable particles. The detector response is also included in our simulations with the use of Delphes-3.5.0 deFavereau:2013fsa . Since we expect heavy stops which are more likely beyond the reach of the current experiments, we simulate similar analyses reported in Ref.CMS:2025ttk for FCC-hh with 100 TeV Center of Mass (COM) energy, by employing FCChh_II.tcl card within Delphes, which is generated suitably with the official design of FCC-hh. We discuss our results at this step in terms of the statistical significance for the signal events, which is calculated as follows Cowan:2010js ; Kumar:2015tna :
$\mathcal{Z}=\sqrt{2\left[(S+B)\,\ln\left(1+\dfrac{S}{B}\right)-S\right]},$ (10)
where $S$ represents the number of events for the signal, while $B$ for the total background processes.
4 Fundamental Parameter Space of YU
In this section, we discuss the impact from the NH terms in the fundamental parameter space of YU and the fine-tuning measure. In the first part, we will give a comparison between results from two independent sets of scans. One set called “Holomorphic” does not take the NH contributions into account, while the other (denoted by NH) assumes the presence of these terms. Both sets are optimized to have statistically well-distributed YU solutions. After the comparison, we will discuss the differences in terms of the NH threshold corrections defined in Eq.(3). We will also include briefly the Higgsino-like LSP solutions and their DM implications.
4.1 Comparison between Holomorphic and NH Cases
Figure 3: The results for YU in correlation with $m_{0}$ (top) and $\tan\beta$ (bottom) for the holomorphic (left) and NH (right) cases. All the solutions are compatible with REWSB and LSP neutralino condition. The green points satisfy the mass bounds and constraints from rare $B$-meson decays. The red points form a subset of green, and they represent the relic density solutions compatible with the Planck measurements within $5\sigma$. The horizontal dashed lines indicate the solutions with $R_{tb\tau}=1.1$, and the points below these lines are identified as the YU solutions.
Figure 4: The plots in the $R_{tb\tau}-m_{\tilde{g}}$ (top) and $R_{tb\tau}-m_{\tilde{t}_{1}}$ (bottom) planes for the holomorphic (left) and NH (right) cases. The color coding and the dashed line represent the solutions as described in Figure 3. The gluino mass bounds are not applied to the top planes, while these bounds are represented with vertical dashed lines.
Figure 5: The plots in the $R_{tb\tau}-\mu$ in the holomorphic (left) and NH (right) cases. The colors and the dashed lines represent the solutions as described in Figure 3
We display the YU solutions parametrized by $R_{tb\tau}$ in correlation with the universal scalar mass, $m_{0}$ (top) and $\tan\beta$ (bottom) in Figure 3. The left panels show the results for the holomorphic case, while the right panels include the NH contributions. These two data sets are generated independently to accommodate possible best YU solutions in both cases. All the solutions are compatible with REWSB and LSP neutralino condition. The green points satisfy the mass bounds and constraints from rare $B$-meson decays. The red points form a subset of green, and they represent the relic density solutions compatible with the Planck measurements within $5\sigma$. The horizontal dashed lines indicate the solutions with $R_{tb\tau}=1.1$, and the points below these lines are identified as the YU solutions. As mentioned before, the YU condition brings a strong impact in the low scale implications which is also seen in the $R_{tb\tau}-m_{0}$ plane in the top-left. In the holomorphic case, the YU solutions can be realized at heavy scales of $m_{0}$ ($\gtrsim 10$ TeV), and seeking for the correct relic density of LSP neutralino (red) shifts it to much heavier mass scales. Consequently, these results lead to heavy mass spectra for the scalars at the low scale. On the other hand, when the NH contributions are included one can accommodate the YU solutions in the region where $m_{0}$ is as low as about 1 TeV, as seen from the top-right plane. The strong impact of YU can also be observed in the correlation between $R_{tb\tau}$ and $\tan\beta$ shown in the bottom planes. The YU solutions require a very narrow range ($45\lesssim\tan\beta\lesssim 50$), and if one requires a perfect YU solution ($R_{tb\tau}=1$), it can be realized specifically when $\tan\beta\simeq 47$. This specific value for $\tan\beta$ is determined mostly by the mass ratio of the top and bottom quarks. The NH contributions loose the strong correlation between YU and $\tan\beta$ as seen from the bottom-right plane. YU can be accommodated in a wider range of $\tan\beta$ as $43\lesssim\tan\beta\lesssim 60$. Moreover, the perfect YU solutions can also be realized abundantly in the range $45\lesssim\tan\beta\lesssim 52$.
Impact of the results discussed in the top planes of Figure 3 can immediately be seen in the mass spectrum compatible with the YU condition. Figure 4 displays our results in the $R_{tb\tau}-m_{\tilde{g}}$ (top) and $R_{tb\tau}-m_{\tilde{t}_{1}}$ (bottom) planes for the holomorphic (left) and NH (right) cases. The color coding and the horizontal dashed lines represent the solutions as described in Figure 3. The gluino mass bounds are not applied to the top planes, while these bounds are represented with vertical dashed lines. As discussed in Section 2, the gluino should be light to keep the positive threshold corrections small in the holomorphic case, and the $R_{tb\tau}-m_{\tilde{g}}$ plane in the top-left shows directly its consequence. The solutions in the regions between the two vertical dashed lines are allowed only when gluino is NLSP. The results show that the consistent gluino masses can barely fit with the YU condition (green), and these solutions cannot be compatible with the desired relic density for LSP neutralino to be accounted for the DM observations. YU becomes quite worse in the region of heavy gluino ($R_{tb\tau}\gtrsim 1.2$). The positive threshold corrections from the first term in Eq.(1) are so large in this region that the Yukawa couplings cannot unify. On the other hand, these large positive contributions can be compensated with the negative NH contributions such that the YU can be accommodated with heavy gluino solutions as shown in the top-right plane of Figure 4. Similarly, in the holomorphic case, the stop can only be as light as about 2.2 TeV, while with the NH contributions its mass can be lowered to about 1.5 TeV as shown in the bottom plane. Even though the lowest mass scales for stops are close in both cases, the main impact can be seen in the DM implications. In the holomorphic case, the correct relic density of LSP neutralino (red) can be obtained when the stop weighs about 6 TeV, while the NH case abundantly fit the light stop solutions with the suitable DM relic density.
We conclude the comparison between the holomorphic and NH cases with the allowed ranges for $\mu$ in Figure 5 with plots in the $R_{tb\tau}-\mu$ in the holomorphic (left) and NH (right) cases. The colors and the dashed lines represent the solutions as described in Figure 3. As discussed before, YU favors large $\mu$ values in the holomorphic case, and the left panel can accommodate the YU solutions in the regions with $\mu\gtrsim 4$ TeV. On the other hand, the desired negative contributions to $y_{b}$ can be provided with $\mu^{\prime}$ and $A_{b}^{\prime}$ as given in Eq.(3). With these NH contributions, the $\mu$-term can be even as low as about 100 GeV as shown in the right plane of Figure 5. Indeed, the NH contributions puts an upper bound on $\mu$ such that it cannot be larger than about 10 TeV. Beyond these values, $y_{b}$ starts receiving excessively negative contributions which also break YU.
4.2 YU, Low Fine-Tuning and Higgsino-like LSP
Figure 6: The threshold corrections to $y_{b}$ in the $\delta_{y_{b}}^{H}-\delta_{y_{b}}^{NH}$, $\delta_{y_{b}}^{NH}-A_{b}^{\prime}/m_{\tilde{b}}$, $R_{tb\tau}-\delta_{y_{b}}$ and $\delta_{y_{b}}-\mu$ planes. The color coding in the bottom-left plane is the same as in Figure 3. In the other planes, the green points show the solutions consistent with the mass bounds and constraints from rare $B$-meson decays. The orange points form a subset of green which are compatible with the YU condition. Red solutions on top of orange represent the YU solutions with the relic density allowed by the Planck measurements within $5\sigma$.
Figure 7: The impact of the NH terms in the fine-tuning and DM implications. The color coding in the top-left plane is the same as in Figure 3, while the other planes follow the color coding as described for the top planes in Figure 6. The horizontal dashed lines in the $\Omega h^{2}-m_{\tilde{H}}$ represent the current measurement of Planck satellite within $5\sigma$. The solid (dashed) curves in the bottom-right plane represents the current (future projected) sensitivity of the XENON experiment XENON:2020kmp . The bottom planes display only the solutions in which the Higgsinos form DM by $50\%$ and more.
As is seen above, the NH contributions can significantly widen the parameter space of YU. This is a direct impact from the NH terms through their threshold corrections to $y_{b}$ given in Eq.(3). Indeed, they can completely take over in such corrections such that the YU solutions can be realized almost independently from the holomorphic contributions. We display these corrections in the $\delta_{y_{b}}^{H}-\delta_{y_{b}}^{NH}$, $\delta_{y_{b}}^{NH}-A_{b}^{\prime}/m_{\tilde{b}}$, $R_{tb\tau}-\delta_{y_{b}}$ and $\delta_{y_{b}}-\mu$ planes in Figure 6. The color coding in the bottom-left plane is the same as in Figure 3. In the other planes, the green points show the solutions consistent with the mass bounds and constraints from rare $B$-meson decays. The orange points form a subset of green which are compatible with the YU condition. Red solutions on top of orange represent the YU solutions with the relic density allowed by the Planck measurements within $5\sigma$. The top-left plane compares the holomorphic and NH contributions based on Eqs.(1 and 3). As is seen from the results in this plane, the holomorphic contributions can be positive in most of the YU solutions (orange) and even can be as large as about 2. Such large positive contributions are compensated with large negative NH contributions ($\sim-2.5$). The need for such large negative NH contributions require negative $A_{b}^{\prime}$ as seen in the $\delta_{y_{b}}^{NH}-A_{b}^{\prime}/m_{\tilde{b}}$ plane. The magnitude of $A_{b}^{\prime}$ is about three times as much as the sbottom mass.
The bottom planes of Figure 6 display the total threshold corrections to $y_{b}$ ($\delta y_{b}$) to accommodate the YU solutions. The $R_{tb\tau}-\delta_{y_{b}}$ shows that the total corrections to $y_{b}$ mostly happen to be negative in the range $1\lesssim\delta y_{b}\lesssim 0$ (green under the horizontal dashed line). The relic density constraint (red) shrinks this region as $0.5\lesssim\delta y_{b}\lesssim 0$. Finally we also display the possible ranges for $\mu$-term which make the discussions of fine-tuning and Higgsino-like LSP relevant to our results. The YU solutions (orange) in the $\delta_{y_{b}}-\mu$ can be compatible with the low values of $\mu$ ($\gtrsim 100$ GeV), while it can also be as large as about 10 TeV.
We convert the values of $\mu$ to the fine-tuning measure by following the definitions in Eq.(4 and 5) in Figure 7. The color coding in the top-left plane is the same as in Figure 3, while the other planes follow the color coding as described for the top planes of Figure 6. The horizontal dashed lines in the $\Omega h^{2}-m_{\tilde{H}}$ represent the current measurement of Planck satellite within $5\sigma$. The solid (dashed) curves in the bottom-right plane represents the current (future projected) sensitivity of the XENON experiment XENON:2020kmp . The bottom planes display only the solutions in which the Higgsinos form DM by $50\%$ and more. As is seen from the top-left plane, the NH contributions can significantly improve the fine-tuning results. The YU solutions can be realized with $\Delta_{{\rm EW}}$ as low as about 100, while the relic density constraint (red) can lift the lower bound on $\Delta_{{\rm EW}}$ to about 400. As discussed before, the Higgsino mass does not have to be light for such solutions because of the contributions from $\mu^{\prime}$, and the $m_{\tilde{H}}-\Delta_{{\rm EW}}$ plane shows that the Higgsinos can be as heavy as about 1 TeV in this region. However, despite the significant improvement in $\Delta_{{\rm EW}}$, the Higgsinos receive a strong impact from the Planck measurements and direct detection of DM experiments, when they form LSP. The $\Omega h^{2}-m_{\tilde{H}}$ plane shows that the correct relic density can be satisfied for the Higgsino masses as $0.7\lesssim m_{\tilde{H}}\lesssim 1.5$ TeV, when they participate in LSP composition more than $50\%$. The lighter Higgsino masses lead to low relic density, and these solutions can still take part in DM observations, but they should be accounted for the testable implications partially. One can reflect their partial contributions in DM observations by rescaling their scattering cross-section as Belanger:2015vwa :
$\xi=\left\{\begin{array}[]{ll}1&{\rm for\penalty 10000\ }0.114\leq\Omega h^{2}\leq 0.126\penalty 10000\ ,\\ \dfrac{\Omega h^{2}}{0.12}&{\rm for}\penalty 10000\ \Omega h^{2}<0.114\penalty 10000\ .\end{array}\right$. (11)
Even with such a rescaling, the Higgsino LSP can still be excluded by the current analyses of the direct detection experiments as shown in the bottom-right plane of Figure 7. We display only the solutions for which the LSP neutralino is formed by the Higgsinos more than $50\%$. Despite their rescaled scattering cross-section, the current sensitivity of the XENON experiment (solid curve) is enough to exclude such solutions. The solutions can be compatible with these results only when $m_{\tilde{\chi}_{1}^{0}}\gtrsim 100$ GeV, but these solutions yield quite negligible relic density ($\Omega h^{2}\lesssim 10^{-3}$) such that they cannot contribute to the DM observations.
In conclusion, the DM observations cannot be accommodated when the Higgsinos form the LSP neutralino by $50\%$ or more. A similar discussion can hold also for the Wino-like LSP neutralino. It usually leads to low relic density when it is lighter than about 1 TeV ATLAS:2024qmx , and the current experiments can exclude it up to about 2.8 TeV Fan:2013faa ; MAGIC:2016xys ; HESS:2016mib . Thus, the DM observations in our model can be consistently realized, when the LSP neutralino is formed mostly by Bino.
5 Probing YU through Compressed Spectra: Stop on Neutralino
The discussion in the previous section can be concluded that the NH contributions drastically change the fundamental parameter space such that one can consider re-analysing its low scale implications. In this context, we continue our discussion with the mass scales of supersymmetric particles and their possible collider probes. In this section, we first discuss the strongly interacting sparticles and their status with the bounds on their mass scales from the current analyses. Since the LHC experiments collide the protons, the experimental analyses are expected to be very sensitive to these sparticles. Among them, we discuss the stop separately, since it can be as light as about 1.5 TeV consistent with all the constraints in our results. At the end of this section, we also include detailed collider analyses which can potentially probe stops.
5.1 SUSY Mass Spectrum
Figure 8: The mass spectrum of SUSY particles in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{g}}$, $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{q}_{L}}$, $m_{\tilde{q}_{L}}-m_{\tilde{g}}$ and $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{b}_{1}}$ planes. The color coding is the same as in the top planes of Figure 6. The gluino mass bound is not applied to the results in the left panels. Diagonal lines show the degenerate solutions in masses of the plotted particles. The solid and dashed curves represent the current exclusion bounds as a function of mass from the experimental analyses ATLAS:2020syg ; ATLAS:2021yij .
Figure 8 displays our results for the mass spectra of the supersymmetric particles in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{g}}$, $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{q}_{L}}$, $m_{\tilde{q}_{L}}-m_{\tilde{g}}$ and $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{b}_{1}}$ planes. The color coding is the same as in the top planes of Figure 6. The gluino mass bound is not applied to the results in the left panels. Diagonal lines show the degenerate solutions in masses of the plotted particles. The solid and dashed curves represent the current exclusion bounds as a function of mass from the experimental analyses ATLAS:2020syg ; ATLAS:2021yij as indicated in legends.
The top-left plane represents our results for the gluino mass in correlation with the neutralino mass. The mass degeneracy between the gluino and LSP neutralino (the orange and red solutions around the diagonal line) can be realized in a gluino mass range $0.7\lesssim m_{\tilde{g}}\lesssim 1.5$ TeV, and these solutions can be tested through the gluino pair production which proceeds with gluino decaying into a pair of quarks along with the LSP neutralino (dashed curve) or chargino (solid curve) ATLAS:2020syg . These analyses can exclude the gluino in this region up to about 1 TeV in its neutralino-decays, and even a stronger exclusion can be realized up to about 1.4 TeV in its chargino-decays. These results leave a very small window for the NLSP gluino solutions which are more likely tested very soon. In this region, the correct relic density for LSP neutralino (red points) can be realized through gluino-neutralino coannihilation scenario; thus the collider analyses will also test the DM implications as well. The NH contributions, as mentioned before, allow the heavy gluino solutions, and our results show that YU can be realized when the gluino mass lies from about 2.2 TeV to 10 TeV, as well. The current collider experiments are projected to probe these solutions up to about 2.5 TeV, while FCC can significantly upgrade the sensitivity in the gluino analyses up to about 6 TeV Altin:2019veq .
The squarks from the first two families also take important part in gluino analyses, and the current results can exclude them up to about 1.5 TeV through their decays into LSP neutralino. On the other hand, YU leads to rather heavy spectra for such squarks that they cannot be lighter than about 5 TeV. As is seen from the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{q}_{L}}$ plane, such high mass scales are quite far from the current reach of the experimental analyses (solid curve) ATLAS:2020syg . The results from these analyses can be converted into the results shown in the $m_{\tilde{q}_{L}}-m_{\tilde{g}}$ plane. As is seen, almost all the YU solutions (orange and red) accumulated in the region which is consistent with the current experimental results. Note that, the solutions in the light gluino region ($m_{\tilde{g}}\lesssim 1.4$ TeV) are also shown in color in this plane. This is because the gluino mass bound is not applied, which excludes almost all the colourful solutions in this region.
The bottom-right plane also shows our result for the sbottom mass scales compatible with YU. After the NH contributions, one can realize sbottom as light as about 1.5 TeV. A small portion of these solutions are already excluded by the current analyses which perform in the region with $m_{\tilde{\chi}_{1}^{0}}\lesssim 800$ GeV ATLAS:2021yij . One of the challenges in sbottom analyses is to have final states with $b$-quarks, which can be identified with about 70% efficiency at the detectors ATLAS:2022miz . Thus, when the LSP neutralino weighs heavier than about 800 GeV, the solutions with $m_{\tilde{b}_{1}}\simeq 1.5$ TeV can be consistent with such analyses. As mentioned before, such light sbottom solutions cannot be involved in the YU compatible spectra in the absence of NH contributions. However, our results show that the sbottom analyses can also provide a probe for YU, when the NH terms are included.
5.2 Stop On Neutralino
Figure 9: The Stop and neutralino masses in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ and $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ planes, where $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})\equiv m_{\tilde{t}_{1}}-m_{\tilde{\chi}_{1}^{0}}$. The color coding is the same as in the top planes of Figure 6. The curves represent the exclusion from the experimental analyses CMS:2021beq ; CMS:2025ttk .
We consider the implications for stops separately from the other sparticles in this section, since it is also of interest in detector analyses. Even though these analyses can currently probe the stops up to about 1.2 TeV CMS:2021beq ; CMS:2025ttk , the main challenge in stop analyses is that possible signal processes yield very similar final states as those in the main background formed by the top-pair production. These similarities are not limited to the final state particles, but they include also the transverse momenta, missing energy, and angular distributions etc. Thus, suppression on the background processes also kills the signal events significantly Cici:2016oqr . In order to realize statistically meaningful number of signal events one needs to consider rather large mass differences between stops and LSP neutralino. In this way, the signal events can have considerable statistical significance enhanced with the integrated luminosity.
Before proceeding into detailed detector analyses, we present our results for mass scales for stop in Figure 9 in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ and $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ planes, where $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})\equiv m_{\tilde{t}_{1}}-m_{\tilde{\chi}_{1}^{0}}$. The color coding is the same as in the top planes of Figure 6. The curves represent the exclusion from the experimental analyses CMS:2021beq ; CMS:2025ttk . The YU solutions (orange) around the exclusion curve can yield stops as light as about 1.5 TeV which are slightly beyond the exclusion limit as seen in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ plane. The studies have projected HL-LHC sensitivity to probe such stop solutions up to about 2.2 TeV Baer:2023uwo ; Baer:2025zqt , and the FCC-hh experiments will more likely improve the results up to about 3 TeV Altin:2020qmu .
The interesting observation in these results can be realized in the region where the stop and LSP neutralino are nearly degenerate in mass. Such solutions are not present in the holomorphic case Hussain:2025bxp . Even though such solutions can escape from the sensitivity of the analyses mentioned above, they can be subjected to the analyses over compressed spectra in which the stop does not weigh much heavier than LSP neutralino. Recently revealed results from these analyses CMS:2025ttk are represented with the dashed and solid curves in the $\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})-m_{\tilde{t}_{1}}$ plane. However, these analyses can probe the stop up to about 700 GeV yet, while the YU solutions appear in the region with $m_{\tilde{t}_{1}}\gtrsim 1.4$ TeV. These solutions are quite beyond the sensitivities of the current collider experiments, but they can be tested in future experiments, which are proposed to operate with much higher center of mass energies such as FCC-hh.
Figure 10: The Significance of stop-pair production in correlation with the mass difference between the stop and LSP neutralino at 100 TeV COM with integrated luminosity ($\mathcal{L}$) is 10 fb-1 (left) and 139 fb-1 (right). In calculation of significance, only the top-quark pair production is considered in the background events such that each stop decays into the LSP neutralino associated with top-quark. The color coding is the same as in the top planes of Figure 6. The horizontal dashed lines represent the significance equal to 1, 2 and 3 from bottom to top, respectively.
Figure 10 displays the significance for the events diagrammatically represented in Figure 2. The significance calculated as given in Eq.(10) projected to the FCC-hh experiments of 100 TeV COM. The color coding is the same as in the top planes of Figure 6. The horizontal dashed lines represent the significance equal to 1, 2 and 3 from bottom to top, respectively. The left plane shows the significance for $\mathcal{L}=10$ fb-1. The YU solutions (orange) can yield stop-pair production events with $\mathcal{Z}\simeq 1$, which can be probed in FCC experiments up to about $68\%$ CL. If one also requires these solutions to be compatible with the Planck measurements on relic abundance of LSP neutralino (red), the maximum significance can be realized at about 0.5. The right panel shows that these results can slightly be improved when the integrated luminosity reaches to about 139 fb-1 as $\mathcal{Z}\simeq 1.2$ for the YU solutions (orange) and $\mathcal{Z}\simeq 0.8$ when the YU solutions are consistent with the DM observations (red).
Point 1 Point 2 Point 3 Point 4
$m_{0}$ $6218.3$ $7210.5$ $7434.1$ $7729.8$
$M_{1}$ $1372.5$ $3006.2$ $1042.1$ $3215.7$
$M_{2}$ $747.3$ $4625.3$ $564.7$ $4944.8$
$M_{3}$ $2310.2$ $577.6$ $1758.3$ $622.0$
$m_{H_{d}}$ $9614.7$ $7805.3$ $12490.6$ $8783.6$
$m_{H_{u}}$ $7423.3$ $4890.8$ $8083.1$ $5516.9$
$A_{0}/m_{0}$ 1.57-1.57 1.71-1.71 1.25-1.25 1.75-1.75
$A_{0}^{\prime}/m_{0}$ 1.57-1.57 1.71-1.71 1.25-1.25 1.75-1.75
$\tan\beta$ $46.9$ $47.0$ $49.1$ $43.6$
$\mu$ $3992.4$ $7037.5$ $3337.1$ $7364.3$
$\mu^{\prime}$ 1792.4-1792.4 3546.1-3546.1 1317.8-1317.8 3861.0-3861.0
$\Delta_{{\rm EW}}$ $3.83\times 10^{3}$ $1.19\times 10^{4}$ $2.68\times 10^{3}$ $1.30\times 10^{4}$
$m_{h}$ $126.8$ $126.7$ $125.9$ $126.8$
$m_{H}$ $4279.5$ $4230.1$ $6626.0$ $4610.4$
$m_{A}$ $4279.5$ $4230.1$ $6626.0$ $4610.4$
$m_{H^{\pm}}$ $4283.4$ $4234.0$ $6630.0$ $4613.7$
$m_{\tilde{\chi}_{1}^{0}}$$m_{\tilde{\chi}_{2}^{0}}$ $\bf\color[rgb]{1,0,0}619.2$} , $\bf\color[rgb]{1,0,0}649.0$} $\bf\color[rgb]{1,0,0}1409.4$} , $3466.6$ $473.2$ , $502.0$ $\bf\color[rgb]{1,0,0}1509.3$} , $3484.8$
$m_{\tilde{\chi}_{3}^{0}}$$m_{\tilde{\chi}_{4}^{0}}$ $2223.7$ , $2224.7$ $3473.6$ , $3966.6$ $2039.8$ , $2040.9$ $3489.5$ , $4234.0$
$m_{\tilde{\chi}_{1}^{\pm}}$$m_{\tilde{\chi}_{2}^{\pm}}$ $\bf\color[rgb]{1,0,0}649.1$} , $2225.7$ $3466.9$ , $3966.6$ $502.1$ , $2041.9$ $3485.0$ , $4234.0$
$m_{\tilde{g}}$ $5063.0$ $\bf\color[rgb]{1,0,0}1509.8$} $4028.3$ $1615.2$
$m_{\tilde{u}_{1}}$$m_{\tilde{u}_{2}}$ $7285.7$ , $7401.3$ $7178.1$ , $7810.9$ $7802.9$ , $8052.1$ $7697.9$ , $8366.9$
$m_{\tilde{t}_{1}}$$m_{\tilde{t}_{2}}$ $2972.8$ , $3243.4$ $2286.7$ , $3785.4$ $3109.5$ , $3730.9$ $\bf\color[rgb]{1,0,0}1713.4$} , $3720.6$
$m_{\tilde{d}_{1}}$$m_{\tilde{d}_{2}}$ $7401.7$ , $7426.1$ $7302.4$ , $7811.4$ $8052.5$ , $8109.9$ $7827.3$ , $8367.3$
$m_{\tilde{b}_{1}}$$m_{\tilde{b}_{2}}$ $2944.4$ , $3136.9$ $2438.6$ , $3773.7$ $\bf\color[rgb]{1,0,0}1435.9$} , $3128.1$ $2218.1$ , $3708.7$
$m_{\tilde{\nu}_{e}}$$m_{\tilde{\nu}_{\tau}}$ $6138.7$ , $4353.2$ $7703.5$ , $5868.0$ $7259.6$ , $5101.6$ $8254.1$ , $6499.0$
$m_{\tilde{e}_{1}}$$m_{\tilde{e}_{2}}$ $6139.6$ , $6399.8$ $7434.4$ , $7704.3$ $7260.5$ , $7756.0$ $7965.3$ , $8254.9$
$m_{\tilde{\tau}_{1}}$$m_{\tilde{\tau}_{2}}$ $1863.1$ , $4354.4$ $2307.4$ , $5868.9$ $2580.2$ , $5102.7$ $3397.0$ , $6499.9$
$\Omega h^{2}$ $0.121$ $0.124$ $0.121$ $0.126$
$\sigma_{{\rm SI}}$ $1.57\times 10^{-12}$ $1.35\times 10^{-12}$ $1.46\times 10^{-12}$ $1.73\times 10^{-12}$
$\sigma_{{\rm SD}}$ $7.56\times 10^{-9}$ $1.14\times 10^{-9}$ $1.02\times 10^{-8}$ $1.35\times 10^{-9}$
$\sigma({\rm Signal})_{14}$ $1.81\times 10^{-9}$ $8.85\times 10^{-9}$ $5.68\times 10^{-12}$ $7.96\times 10^{-5}$
$\sigma({\rm Signal})_{100}$ $1.77\times 10^{-4}$ $1.06\times 10^{-4}$ $8.60\times 10^{-7}$ $\bf\color[rgb]{1,0,0}1.90\times 10^{-1}$}
$\mathcal{Z}({\rm Signal})_{100}^{10}$ $0.0$ $0.0$ $0.0$ $0.5$
$R_{tb\tau}$ $\bf\color[rgb]{1,0,0}1.01$} $1.09$ $1.07$ $1.09$
Table 1: Benchmark scenarios summarizing our findings. In selection the solutions are required to be consistent with all the constraints. All the masses are given in GeV and cross-sections in pb. The subscripts in the cross-section expressions indicate the COM of the collider experiments. The colored values indicate the emphasized features of selected benchmark points.
The significance is calculated at this level very roughly by assuming only the top-quark pair production forming the SM background. In addition, the kinematic variables are not considered. Before proceeding into the detailed analyses and improved significance, we first summarize our findings with four benchmark points in Table 1. We select these points by requiring to be consistent with all the constraints. The masses are given in GeV, while the cross-sections in pb. The subscripts in the cross-sections show the COM of the collider experiments. The statistical significance of the signal events is given for the collisions with 100 COM (shown with the subscript of $\mathcal{Z}$) and 10 fb-1 integrated luminosity (shown with superscript of $\mathcal{Z}$). The colored values indicate the emphasized features of selected benchmark points. Point 1 represents the solutions with almost perfect YU. The spectra of such solutions involve typically heavy sparticles, and the Planck measurements on relic density of LSP neutralino can be realized through chargino-neutralino coannihilation scenario. Points 2 depicts NLSP gluino solutions which are consistent with the current analyses, and it exemplifies also gluino-neutralino coannihilation scenario. Point 3 displays the possible lightest sbottom which is expected to be tested soon in collider analyses. Point 4 represents stop solutions with possible maximum significance consistent with all the constraints employed in our analyses. The mass difference between the stop and LSP neutralino is about 200 GeV, and it can yield roughly $\mathcal{Z}\simeq 0.5$.
5.3 FCC Probe of Compressed Stop-Neutralino
The benchmark points given in Table 1 can be tested in several analyses as discussed above, but they do not mimic in stop searches except Point 4. In this section we consider Point 4 as a benchmark scenario, and perform detailed analyses to explore possible traces in the future collider experiments. In our work we simulate similar analyses reported in Ref. CMS:2025ttk as briefly described in Section 3.
Process $\sigma_{\text{LO}}$ [14 TeV] (pb) K-Factor [14 TeV] $\sigma_{\text{LO}}$ [100 TeV] (pb) K-Factor [100 TeV]
Signal
($pp\to\tilde{t}\tilde{t}^{*}$), $(\tilde{t}\to t\tilde{\chi}_{1}^{0},\bar{\tilde{t}}\to\bar{t}\tilde{\chi}_{1}^{0})$ $7.96\times 10^{-5}$ - $1.90\times 10^{-1}$ -
Backgrounds
$pp\rightarrow t\bar{t}$ $5.056\times 10^{2}$ $1.67$ $2.469\times 10^{4}$ $1.5$
$pp\rightarrow tW^{\pm}$ $5.501\times 10^{1}$ $1.27$ $2.268\times 10^{3}$ $1.5$
$pp\rightarrow t\bar{t}V(V=Z,W^{\pm})$ $9.36\times 10^{-1}$ $1.3$ $4.953\times 10^{1}$ $1.5$
$pp\rightarrow VV(V=Z,W^{\pm})$ $9.928\times 10^{1}$ $1.35$ $1.022\times 10^{3}$ $1.5$
Table 2: The signal and relevant background process at the leading order (LO) and K-factors.
Table 2 lists the cross-sections for the signal and relevant background processes at 14 TeV and 100 TeV COM. These values are calculated at the leading order (LO), but the next to LO (NLO) contributions are also included by multiplying with the K-factors Kidonakis:2023juy ; Frixione:2008yi ; Garzelli:2012bn ; Grazzini:2019jkl . Note that an average value is employed for the K-factor when 100 TeV COM is considered Mangano:2016jyj . These K-factors are applied only to the background processes. Before considering the relevant kinematical variables, we first apply the following conditions on the events:
$n_{l}=1,\hskip 8.5359ptn_{j}\geq 4,\hskip 8.5359ptn_{b}\geq 2,$ (12)
where $n_{l,j,b}$ denote the numbers of leptons, jets and $b$-quarks in the final states, respectively. These conditions can provide suitable pre-selection rules when the semi-leptonic final states are considered as shown in Figure 2. The further cuts on the kinematic variables are applied on these pre-selection rules. Some of the kinematic variables which can distinguish the signal events from those forming the background are shown in Figure 11. The event numbers for the signal and background processes are normalized to 1. $E_{T}^{{\rm Miss}}$ and $H_{T}$ represent the missing and hadronic transverse energies, respectively. $R_{{\rm ISR}}$ denotes the projection of missing transverse energy into the direction of jets emerging from initial state radiation (ISR). Finally $\Delta R_{{\rm min}}$ is the minimum of the angular distance between the variables denoted in parenthesis. The distribution for the total background is represented with the red curve, while the blue curve shows for the signal processes.
Figure 11: The kinematic variables which can distinguish the signal events from those forming the background. The event numbers for the signal and background processes are normalized to 1. $E_{T}^{{\rm Miss}}$ and $H_{T}$ represent the missing and hadronic transverse energies, respectively. $R_{{\rm ISR}}$ denotes the projection of missing transverse energy into the direction of jets emerging from initial state radiation (ISR). Finally $\Delta R_{{\rm min}}$ is the minimum of the angular distance between the variables denoted in parenthesis. The distribution for the total background is represented with the red curve, while the blue curve shows for the signal processes.
$E_{T}^{{\rm Miss}}$ is typically one of the key variable since the signal processes usually have more missing energy due to the presence of LSP neutralinos. However, this is true when the mass difference between the stop and LSP neutralino is large enough, which is not applicable in our case. The mass difference in our benchmark point is about 200 GeV, and it does not yield a strong impact on the missing energy in compared to the masses of the stop and LSP neutralino. As is seen in the plane plotting $E_{T}^{{\rm Miss}}$, the signal and background events lead to a peak at a similar scale, at which the missing energy is formed mostly by neutrinos. However, beyond this scale, the missing energy distribution exhibits a sharp decrease for the background events, while its change is relatively slow for the signal. It is observed because of the small but non-zero transverse momenta of the LSP neutralinos in the signal processes. Another variable can be formed by considering the ratio of the hadronic transverse energy to the missing energy as shown in the top-right plane. Even though this ratio yield similar peaks as observed in the missing energy distribution, it decrease much faster for the signal than the background processes.
Another important variable can arise from the ISR jets. In the compressed spectrum, most of the energy is consumed by the LSP neutralino mass which softens the top quark. In this case, the ISR jet and $E_{T}^{{\rm Miss}}$ are in the opposite directions. The projection of $E_{T}^{{\rm Miss}}$ onto the ISR jet momentum is expected to be maximum at about ratio of the LSP neutralino mass to stop mass ($\sim 1$ in our case), while it is observed small and wider for the background processes ATLAS:2017drc . This discussion for $R_{{\rm ISR}}$ can be observed in the bottom-left plane. Finally we also display the minimum angular distance between $E_{T}^{{\rm Miss}}$ and the jets in the final states. The distribution given in the bottom-right plane can provide an improvement to isolate the signal events by imposing a cut on the minimum angular distance between jets and $E_{T}^{{\rm Miss}}$, since the signal process exhibits a peak at a slightly lower rate of $\Delta R_{{\rm min}}(j_{i},E_{T}^{{\rm Miss}})$.
$E_{T}^{{\rm Miss}}$ $H_{T}/E_{T}^{{\rm Miss}}$ $R_{{\rm ISR}}$ $\Delta R_{{\rm min}}(j_{i},E_{T}^{{\rm Miss}})$
Set 1 $\geq$ 665.0 GeV - - -
Set 2 $\geq$ 665.0 GeV $\leq$ 3.33 - -
Set 3 $\geq$ 665.0 GeV $\leq$ 3.33 $\geq$ 0.64 -
Set 4 $\geq$ 665.0 GeV $\leq$ 3.33 $\geq$ 0.64 $\leq$ 5.11
Table 3: The sets of cut flows imposed in our analyses. The cut on each variable is applied subsequently on top of the previous set.
Based on the results discussed with the plots in Figure 11 we impose the cuts on the kinematic variables subsequently as listed in Table 3. The final results are obtained after the fourth set of the cut flow, but we list them as different sets to discuss the effects from each cut in details. Recall that these cuts are imposed after the pre-selection rules are applied. The calculated number of events for FCC-hh of 100 TeV COM are listed in Table 4 after each set. The effectiveness of each cut set on the signal processes is parametrized by $\epsilon_{S}$, which is simply the ratio of number of signal events as
$\epsilon_{S}\equiv\dfrac{{\rm Number\penalty 10000\ of\penalty 10000\ Events\penalty 10000\ (after\penalty 10000\ cut)}}{{\rm Number\penalty 10000\ of\penalty 10000\ Events\penalty 10000\ (pre\text{-}selection)}}\penalty 10000\ ,$ (13)
and the significance $\mathcal{Z}$ is calculated through Eq.(10). After including the relevant background processes enhanced with K-factors, the significance of the pre-selected signal events can be realized to be about 0.5 at about ten times larger integrated luminosity ($\mathcal{L}=100$ fb-1) than that reported in Table 1.
Cut Flow Signal $\bar{t} tt$VttV tWtW VVVV Total Bkg. $\epsilon_{S}$ $\mathcal{Z}$
Pre-selection $2.87\times 10^{4}$ $3.58\times 10^{9}$ $7.18\times 10^{6}$ $3.28\times 10^{8}$ $1.48\times 10^{8}$ $4.06\times 10^{9}$ 1.000 0.450
Set 1 $5.00\times 10^{3}$ $4.69\times 10^{5}$ $1.95\times 10^{4}$ $6.90\times 10^{5}$ $6.38\times 10^{4}$ $1.24\times 10^{6}$ 0.175 4.486
Set 2 $4.72\times 10^{3}$ $3.52\times 10^{5}$ $1.54\times 10^{4}$ $4.45\times 10^{5}$ $5.32\times 10^{4}$ $8.65\times 10^{5}$ 0.165 5.070
Set 3 $4.60\times 10^{3}$ $3.52\times 10^{5}$ $1.45\times 10^{4}$ $4.00\times 10^{5}$ $4.26\times 10^{4}$ $8.09\times 10^{5}$ 0.161 5.111
Set 4 $4.60\times 10^{3}$ $3.52\times 10^{5}$ $1.43\times 10^{4}$ $3.56\times 10^{5}$ $4.26\times 10^{4}$ $7.64\times 10^{5}$ 0.160 5.251
Table 4: The number of events for the signal and background processes realized at FCC-hh experiments of 100 TeV COM and $100$ fb-1 integrated luminosity together with the effectiveness of cut flows and resultant significance for the signal processes.
It can be seen easily from the results given in Table 4 that the cut on the missing energy immediately increases the significance of the signal events by about 10 times, while the other sets can also provide minor improvements. Eventually, only about $16\%$ of the signal events can survive after cuts in Set 4, which corresponds to $\mathcal{Z}\simeq 5.2$. Recall that we assume $\mathcal{L}=100$ fb-1 in calculation of significance, which can be collected on the first days of FCC-hh operation. In this context, one can expect significantly improved results as the integrated luminosity increases, as shown in Figure 12. According to the correlation between the significance and integrated luminosity, the compressed spectra of the stop and LSP neutralino can be probed up to about $3\sigma$ when $\mathcal{L}\simeq 32$ fb-1, and $5\sigma$ probe can be realized for $\mathcal{L}\simeq 97$ fb-1.
Figure 12: The correlation between the significance and integragted luminosity ($\mathcal{L}$). The green dashed line indicates the luminosity value where the significance of the signal events can be realized within $3\sigma$, while the purple dashed line shows its probe within $5\sigma$ at the collider experiments proposed in FCC-$h$.
BDT Improve:
Even though the cut-based analyses can yield a strong probe at the FCC-hh experiments, the results obtained above can be improved further through Boosted Decision Tree (BDT) analyses. We employ the BDT algorithm based on XGBoost (eXtreme Gradient Boost) trained with $10^{5}$ events for each of the signal and background processes generated independently. In addition to the variables described above, the events are transferred to the BDT scan with $m_{CT}$ Tovey:2008ui , $m_{T}\!\left(\ell,E_{T}^{\mathrm{miss}}\right)$ ATLAS:2024cmj , $m_{T2}^{W}$ Bai:2012gs , $m(b,l)$, $m_{T}\left(b_{i},E_{T}^{\mathrm{miss}}\right)$ ATLAS:2017drc . Note that the events are refined only by applying the pre-selection rules without any further cut on the variables.
The expected impact from these variables can be summarized as follows: The number of events reach its peak at about $m_{CT}\approx\Delta m(\tilde{t}_{1},\tilde{\chi}_{1}^{0})$ for the signal, while its distribution is expected to be wider for the background. The transverse mass of lepton and $E_{T}^{{\rm Miss}}$ ($m_{T}\!\left(\ell,E_{T}^{\mathrm{miss}}\right)$) cannot exceed the $W$-boson mass for the background, while it can take larger values for the signal processes, since the LSP neutralino contributes to the missing energy. A cut on $m_{T2}^{W}$ is utilized in such analyses to suppress the decay channels whose visible final state involves a single lepton. The presence of the ISR jet boosts the missing energy, it can provide a useful separation between the signal and background processes in distribution over $m_{T}\left(b_{i},E_{T}^{\mathrm{miss}}\right)$.
In addition, one can consider the Razor variables which are useful when the processes involve invisible particles, especially when they have similar masses. Some of them can be defined as follows CMS:2016igb :
$\begin{array}[]{rl}M_{TR}^{2}&=\dfrac{E_{T}^{\mathrm{Miss}}\left(p_{T1}+p_{T2}\right)-\vec{p}_{T}^{\,\mathrm{Miss}}\cdot\left(\vec{p}_{T1}+\vec{p}_{T2}\right)}{2}\\ M_{R}^{2}&=\left(\left|\vec{p}_{1}\right|+\left|\vec{p}_{2}\right|\right)^{2}-\left(p_{1z}+p_{2z}\right)^{2}\\ R^{2}&=\left(\dfrac{M_{TR}}{M_{R}}\right)^{2}\end{array}$ (14)
If one assigns the top quarks to be megajets, whose transverse momenta are $p_{T1}$ and $p_{T2}$, $R^{2}$ could be useful in distinguishing the signal processes from the background. The boost on $E_{T}^{{\rm Miss}}$ from the ISR jet aligns it with these megajets, which maximize $M_{TR}^{2}$. In this case, one can realize $R^{2}\simeq R^{2}_{{\rm max}}=(m_{\tilde{t}_{1}}-m_{\tilde{\chi}_{1}^{0}})/m_{\tilde{\chi}_{1}^{0}}$. While most of the signal events are expected at $R^{2}\simeq R^{2}_{{\rm max}}$, the distribution of the background events over $R^{2}$ happens to be smoother.
Figure 13: The separation between the signal and background processes by BDT (left) and the best score (right) obtained within our analyses.
We transfer the pre-selected events to BDT algorithm, which is trained together with the variables described above. We obtain the best models with the hyperparameters summarized in Table 5.
Hyperparameter Value Hyperparameter Value
learning_rate $0.0408$ gamma $1.98$
max_depth 66 reg_lambda $0.431$
min_child_weight $5.79$ reg_alpha $3.28$
subsample $0.752$ ROC-AUC 0.933
Table 5: The hyperparameters for the best XGBoost model.
After training, BDT can identify the signal processes over the background quite efficiently as shown in the left panel of Figure 13. The best score in our analyses is obtained as $\tau=0.98$, which is tested over the signal significance displayed in the right plane. With the best score in BDT, we realize a significant improvement that the analyses can probe the signal up to about $3\sigma$ with the integrated luminosity of about $19.4$ fb-1, and $5\sigma$ when $\mathcal{L}\simeq 54.4$ fb-1. The significance of the signal in correlation with the integrated luminosity is given in Figure 14. One can expect from these results that the compressed spectra for the stop and neutralino are more likely to be probed on the first days when FCC-hh starts being operated.
Figure 14: The correlation between the significance and integragted luminosity ($\mathcal{L}$). The green dashed line indicates the luminosity value where the significance of the signal events can be realized within $3\sigma$, while the purple dashed line shows its probe within $5\sigma$ at the collider experiments proposed in FCC-$h$.
6 Conclusion
We explore the implications within the high scale SUSY models based on the gauge symmetry $SU(4)_{C}\times SU(2)_{L}\times SU(2)_{R}$. If its gauge symmetry is broken into the SM gauge group near at $M_{{\rm GUT}}$, one can maintain the gauge coupling unification approximately. In addition, these models also impose YU in their minimal construction. Previous studies have reported rather heavy mass spectra favored by YU. This results is also supported by the correct relic density constraint on the LSP neutralino measured by the Planck satellite. In our work, we also include the NH terms which are mostly ignored in previous studies. These terms significantly change the profile of YU through the threshold contributions to the Yukawa couplings. In this context, we identify the YU solutions compatible with low fine-tuning, while the Higgsino-like LSP can weigh heavier because of the contributions to its mass from $\mu^{\prime}$. Even though this observation can be interesting from naturalness point of view, the Higgsino-like LSP solutions receive strong strike from the DM observations.
The YU condition in the models of 4-2-2 class can be satisfied only when the gluino weighs lighter than about 1.2 TeV, which are more or less excluded by the current collider analyses. On the other hand, the NH contributions make the YU solutions possible in the regions of heavier gluino from about 2.2 to 10 TeV. The collider experiments with high luminosity are projected to probe these solutions up to about 2.5 TeV, while future experiments can improve the sensitivity up to about $m_{\tilde{g}}\lesssim 6$ TeV. The squarks from the first two families are involved in the mass spectra heavier than about 4 TeV, which are beyond the current sensitivity in the collider experiments.
The third family squarks, on the other hand, can be driven to lighter mass scales by the NH terms. Typical YU solutions usually lead to sbottoms in spectra heavier than about 5 TeV, while it can be as light as about 1 TeV when the NH contributions are taken into account. The current analyses can exclude these solutions for $m_{\tilde{b}_{1}}\lesssim 1.5$ TeV. Similarly, stops mass can also be found as light as about 1.5 TeV. These solutions can be excluded up to about 1.2 TeV when the LSP neutralino is lighter than about 700 GeV. We also identify solutions in which the stop and LSP neutralino can be nearly degenerate at $m_{\tilde{t}_{1}}\simeq m_{\tilde{\chi}_{1}^{0}}\gtrsim 1.4$ TeV. Despite these solutions can escape from the common stop analyses, but recent analyses can improve the sensitivity up to about $m_{\tilde{t}_{1}}\lesssim 700$ GeV.
In our analyses, we also require the relic density constraint to be satisfied, and we find the lightest possible stop weigh about 1.7 TeV of about 200 GeV mass difference with the LSP neutralino. Such solutions are clearly beyond the reach of the current analyses, but they can be tested in future experiments such as those proposed in FCC-hh with 100 TeV COM. We discuss possible probes for these solutions by considering the collider analyses projected into FCC-hh. In these analyses, we consider semi-leptonic final states emerging from stop pair-production. With suitable selections described with cuts on the kinematic variables, we realize statistical significance at about $5\sigma$ when the integrated luminosity is increased to about $100$ fb-1. We also perform a BDT analysis which results in such a large significance when the experimental analyses have lower luminosity as $\mathcal{L}\simeq 54$ fb-1. Such luminosity values can be reached on the first days of FCC-hh experiments.
Acknowledgment
The work of BN and CSU is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK) Grant No. MFAG-125F122. The numerical calculations reported in this paper were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).

### 30 Mar 2026 v2

@@
 The detector response is also included in our simulations with the use of Delphes-3.5.0 deFavereau:2013fsa .
-Since we expect heavy stops which are more likely beyond the reach of the current experiments, we simulate similar analyses reported in Ref.CMS:2025ttk for FCC-hh with 100 TeV Center of Mass (COM) energy, by employing FCChh_II.tcl card within Delphes, which is generated suitably with the official design of FCC-hh.
+Since we expect heavy stops which are more likely beyond the reach of the current experiments, we simulate similar analyses reported in Ref.CMS:2025ttk for FCC-$h$ with 100 TeV Center of Mass (COM) energy, by employing FCChh_II.tcl card within Delphes, which is generated suitably with the official design of FCC-$h$.
 We discuss our results at this step in terms of the statistical significance for the signal events, which is calculated as follows Cowan:2010js ; Kumar:2015tna :
@@
 The YU solutions (orange) around the exclusion curve can yield stops as light as about 1.5 TeV which are slightly beyond the exclusion limit as seen in the $m_{\tilde{\chi}_{1}^{0}}-m_{\tilde{t}_{1}}$ plane.
-The studies have projected HL-LHC sensitivity to probe such stop solutions up to about 2.2 TeV Baer:2023uwo ; Baer:2025zqt , and the FCC-hh experiments will more likely improve the results up to about 3 TeV Altin:2020qmu .
+The studies have projected HL-LHC sensitivity to probe such stop solutions up to about 2.2 TeV Baer:2023uwo ; Baer:2025zqt , and the FCC-$h$ experiments will more likely improve the results up to about 3 TeV Altin:2020qmu .
 The interesting observation in these results can be realized in the region where the stop and LSP neutralino are nearly degenerate in mass.
@@
 However, these analyses can probe the stop up to about 700 GeV yet, while the YU solutions appear in the region with $m_{\tilde{t}_{1}}\gtrsim 1.4$ TeV.
-These solutions are quite beyond the sensitivities of the current collider experiments, but they can be tested in future experiments, which are proposed to operate with much higher center of mass energies such as FCC-hh.
+These solutions are quite beyond the sensitivities of the current collider experiments, but they can be tested in future experiments, which are proposed to operate with much higher center of mass energies such as FCC-$h$.
 Figure 10: The Significance of stop-pair production in correlation with the mass difference between the stop and LSP neutralino at 100 TeV COM with integrated luminosity ($\mathcal{L}$) is 10 fb-1 (left) and 139 fb-1 (right).
@@
 Figure 10 displays the significance for the events diagrammatically represented in Figure 2.
-The significance calculated as given in Eq.(10) projected to the FCC-hh experiments of 100 TeV COM.
+The significance calculated as given in Eq.(10) projected to the FCC-$h$ experiments of 100 TeV COM.
 The color coding is the same as in the top planes of Figure 6.
@@
 Recall that these cuts are imposed after the pre-selection rules are applied.
-The calculated number of events for FCC-hh of 100 TeV COM are listed in Table 4 after each set.
+The calculated number of events for FCC-$h$ of 100 TeV COM are listed in Table 4 after each set.
 The effectiveness of each cut set on the signal processes is parametrized by $\epsilon_{S}$, which is simply the ratio of number of signal events as
@@
 Set 4 $4.60\times 10^{3}$ $3.52\times 10^{5}$ $1.43\times 10^{4}$ $3.56\times 10^{5}$ $4.26\times 10^{4}$ $7.64\times 10^{5}$ 0.160 5.251
-Table 4: The number of events for the signal and background processes realized at FCC-hh experiments of 100 TeV COM and $100$ fb-1 integrated luminosity together with the effectiveness of cut flows and resultant significance for the signal processes.
+Table 4: The number of events for the signal and background processes realized at FCC-$h$ experiments of 100 TeV COM and $100$ fb-1 integrated luminosity together with the effectiveness of cut flows and resultant significance for the signal processes.
 It can be seen easily from the results given in Table 4 that the cut on the missing energy immediately increases the significance of the signal events by about 10 times, while the other sets can also provide minor improvements.
 Eventually, only about $16\%$ of the signal events can survive after cuts in Set 4, which corresponds to $\mathcal{Z}\simeq 5.2$.
-Recall that we assume $\mathcal{L}=100$ fb-1 in calculation of significance, which can be collected on the first days of FCC-hh operation.
+Recall that we assume $\mathcal{L}=100$ fb-1 in calculation of significance, which can be collected on the first days of FCC-$h$ operation.
 In this context, one can expect significantly improved results as the integrated luminosity increases, as shown in Figure 12.
@@
 BDT Improve:
-Even though the cut-based analyses can yield a strong probe at the FCC-hh experiments, the results obtained above can be improved further through Boosted Decision Tree (BDT) analyses.
+Even though the cut-based analyses can yield a strong probe at the FCC-$h$ experiments, the results obtained above can be improved further through Boosted Decision Tree (BDT) analyses.
 We employ the BDT algorithm based on XGBoost (eXtreme Gradient Boost) trained with $10^{5}$ events for each of the signal and background processes generated independently.
@@
 The significance of the signal in correlation with the integrated luminosity is given in Figure 14.
-One can expect from these results that the compressed spectra for the stop and neutralino are more likely to be probed on the first days when FCC-hh starts being operated.
+One can expect from these results that the compressed spectra for the stop and neutralino are more likely to be probed on the first days when FCC-$h$ starts being operated.
 Figure 14: The correlation between the significance and integragted luminosity ($\mathcal{L}$).
@@
 In our analyses, we also require the relic density constraint to be satisfied, and we find the lightest possible stop weigh about 1.7 TeV of about 200 GeV mass difference with the LSP neutralino.
-Such solutions are clearly beyond the reach of the current analyses, but they can be tested in future experiments such as those proposed in FCC-hh with 100 TeV COM.
-We discuss possible probes for these solutions by considering the collider analyses projected into FCC-hh.
+Such solutions are clearly beyond the reach of the current analyses, but they can be tested in future experiments such as those proposed in FCC-$h$ with 100 TeV COM.
+We discuss possible probes for these solutions by considering the collider analyses projected into FCC-$h$.
 In these analyses, we consider semi-leptonic final states emerging from stop pair-production.
@@
 We also perform a BDT analysis which results in such a large significance when the experimental analyses have lower luminosity as $\mathcal{L}\simeq 54$ fb-1.
-Such luminosity values can be reached on the first days of FCC-hh experiments.
+Such luminosity values can be reached on the first days of FCC-$h$ experiments.
 Acknowledgment